# Do Not (Always) Look Right: Investigating the Capabilities of Decoder-Based Large Language Models for Sequence Labeling

### 25 Jan 2024 v1

@@ 5.4 Pre-training Small Encoders and Decoders
We inspect whether removing the CM from small CLM-based decoder PLMs (under 100 million parameters) exhibits the same trend Li et al. (2023) observed for high-scale Llama2 models with multiple billion parameters. Furthermore, we verify whether MLM-based pre-training is more appropriate for reaching higher scores on the downstream SL tasks when the number of parameters, pre-training data, pre-training steps, and all other hyperparameters are fixed between encoders and decoders. We randomly initialize small LMs with four encoder or decoder blocks based on RoBERTa-base architecture. We attach a language modeling head to these small LMs and use a randomly initialized embeddings matrix based on RoBERTa-base’s vocabulary. We inherit all other parameters of the RoBERTa-base official configuration, e.g., the hidden size dimensionality of $768$. Initializing this model produces a small LM with $68$ million parameters. We use RoBERTa-base tokenizer with added prefix space to pre-tokenize the “BookCorpus” dataset (Zhu et al., 2015) consisting of $74$ million sentences. We employ this dataset to pre-train our small LMs. “BookCorpus” sentences are grouped to form blocks of size $512$ for either MLM-based pre-training of small encoder or CLM-based pre-training of small decoder LM. MLM probability is set at $0.15$. Small encoder and decoder are trained with AdamW optimizer, a cosine annealing learning rate scheduler, eight gradient accumulation steps, and bfloat16 precision with a batch size of $64$ and a learning rate of $2\mathrm{e}{-4}$. The parameters of AdamW are fixed to $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=1\mathrm{e}{-5},\lambda=0.1$ and we apply gradient clipping to $1.0$. We save the model weights immediately after random initialization and then at five equally spaced intervals throughout each epoch. The training process continues for a pre-determined fixed number of $10$ epochs (roughly 200k steps). We train both models with same seed. Then, we load the weights of each saved checkpoint, replace the LM head with the SL head for the appropriate task, and fine-tune the models for a fixed number of five epochs with a batch size of $16$. We do this for each of the $51$ saved checkpoints with either a pre-trained encoder, a pre-trained decoder with CM kept during fine-tuning, or a pre-trained decoder with CM removed during fine-tuning. We evaluate these variants on validation sets of the four SL datasets. The results are in Figure 2. All results are averages of five fine-tuning runs with different seeds ($seeds=[120,121,122,123,124]$), and we always pick the last model for each seed. The results reveal that removing the CM on this small scale yields no gains. On average, the model without CM pefroms worse than a fine-tuned, pre-trained decoder with CM.
@@
Moreover, small encoders struggle to keep up with decoders until around the twentieth checkpoint (fourth pre-training epoch), when they suddenly start having significantly more success on all SL tasks except trigger classification on ACE05. Training with MLM for over four epochs on the “BookCorpus” dataset drastically hurts performance on ACE05. The score RoBERTa-base achieves on ACE05 (cf. Table 2) is not much higher than the best small encoder. The gains from further MLM pre-training on ACE05 are less accentuated than the overall number of model parameters. This can be corroborated by the fact that RoBERTa-base and RoBERTa-large were both trained for 500k steps (Liu et al., 2019), and RoBERTa-large is drastically better on ACE05.
@@
Figure 2: Validation set micro F1 SL scores of small MLM-based encoder and CLM-based decoder models (pre-trained from scratch under controlled conditions) after fine-tuning on SL tasks training data starting from specific language modeling checkpoint. “Decoder Unmask” means that the model was pre-trained with CLM objective and a CM but fine-tuned without the CM. The shaded area represents the standard deviation.
@@ 6 Conclusion
The landscape of PLMs has undergone a significant transformation, shifting from encoder to decoder dominance in various NLU tasks. This shift is attributed to the scale of decoders and their emergent zero- and few-shot capabilities. While large decoder models like ChatGPT showcase remarkable performance even without fine-tuning, challenges persist in IE-related tasks. “Open” LLMs, such as Llama2 and Mistral, introduce transparency and accessibility, facilitating experimentation with parameter-efficient fine-tuning techniques. Notably, removing the CM during fine-tuning with QLoRA enhances performance on IE tasks modeled as SL, with crucial decisions regarding layer-wise CM removal contributing to substantial improvements.
Future research should delve into refining bidirectional information flow within decoder blocks, addressing task-specific nuances, and predicting the impact of CM removal from specific layer groups without requiring extensive fine-tuning. Comparisons with encoder-based models underscore the significance of architecture and parameter scale in SL tasks. A noticeable gap exists between small-scale encoder and decoder models, where decoder models do not benefit from CM removal at a smaller scale. The ongoing upscaling of decoder models suggests a potential turning point where they might outperform encoders on SL tasks without requiring fine-tuning, leveraging their emerging properties. However, the preferred approach for now involves fine-tuning with a specialized SL head to achieve optimal performance.
@@ Appendix A Appendix
A.1 Instruction Tuning Details
@@
Task (dataset) Instruction example for training Instruction example for evaluation NER (CoNLL03) ### Instruction: please extract named entities and their type from the input sentence, all entity types are in options ### Options: person, location, organization, miscellaneous ### Sentence: " What we have to be extremely careful of is how other countries are going to take Germany ’s lead , " Welsh National Farmers ’ Union ( NFU ) chairman John Lloyd Jones said on BBC radio . ### Response: Germany:location;Welsh National Farmers ’ Union:organization;NFU:organization;John Lloyd Jones:person;BBC radio:organization ### Instruction: please extract named entities and their type from the input sentence, all entity types are in options ### Options: person, location, organization, miscellaneous ### Sentence: " What we have to be extremely careful of is how other countries are going to take Germany ’s lead , " Welsh National Farmers ’ Union ( NFU ) chairman John Lloyd Jones said on BBC radio . ### Response: Aspect Term Extraction and Polarity (ABSA Restaurants) ### Instruction: please extract aspect terms and their polarity from the input sentence, all polarity types are in options ### Options: positive, negative, neutral, conflict ### Sentence: The lobster sandwich is \$ 24 and although it was good it was not nearly enough to warrant that price . ### Response: lobster sandwich:conflict;price:negative ### Instruction: please extract aspect terms and their polarity from the input sentence, all polarity types are in options ### Options: positive, negative, neutral, conflict ### Sentence: The lobster sandwich is \$ 24 and although it was good it was not nearly enough to warrant that price . ### Response: Trigger Classification (ACE05) ### Instruction: please extract events and their types from the input sentence, all event types are in options ### Options: merge organization, start organization, declare bankruptcy, end organization, grant pardon, extradite, execute, impose fine, conduct trial hearing, issue sentence, file appeal, convict, file lawsuit, release on parole, arrest and send to jail, charge and indict, acquit, participate in protest or demonstration, attack, contact via written or telephone communication, meet, start position, elect, end position, nominate, transfer ownership, transfer money, marry, divorce, be born, die, sustain injury, transport ### Sentence: In his previous letter home , Apache pilot Joe Bruhl did n’t tell his family the full details about his first combat mission . ### Response: tell:contact via written or telephone communication;combat:attack ### Instruction: please extract events and their types from the input sentence, all event types are in options ### Options: merge organization, start organization, declare bankruptcy, end organization, grant pardon, extradite, execute, impose fine, conduct trial hearing, issue sentence, file appeal, convict, file lawsuit, release on parole, arrest and send to jail, charge and indict, acquit, participate in protest or demonstration, attack, contact via written or telephone communication, meet, start position, elect, end position, nominate, transfer ownership, transfer money, marry, divorce, be born, die, sustain injury, transport ### Sentence: In his previous letter home , Apache pilot Joe Bruhl did n’t tell his family the full details about his first combat mission . ### Response: Text Chunking (CoNLL03) ### Instruction: please extract chunks and their type from the input sentence, all chunk types are in options ### Options: noun phrase, verb phrase, prepositional phrase, adverb phrase, subordinated clause, adjective phrase, particles, conjunction phrase, interjection, list marker, unlike coordinated phrase ### Sentence: Rare Hendrix song draft sells for almost \$ 17,000 . ### Response: Rare Hendrix song draft:noun phrase;sells:verb phrase;for:prepositional phrase;almost \$ 17,000:noun phrase ### Instruction: please extract chunks and their type from the input sentence, all chunk types are in options ### Options: noun phrase, verb phrase, prepositional phrase, adverb phrase, subordinated clause, adjective phrase, particles, conjunction phrase, interjection, list marker, unlike coordinated phrase ### Sentence: Rare Hendrix song draft sells for almost \$ 17,000 . ### Response:

### 27 Feb 2024 v2

@@ 5.4 Investigating the Effect of Scale
We observed gains upon CM removal on a 7B parameters scale. Our experiments prompt the question of whether similar findings would apply to models of different parameter scales. More specifically, whether CM removal from a small CLM-based decoder would exhibit the same trend, and also whether MLM-based pre-training is more beneficial for success on SL tasks when the number of parameters, pre-training data and steps, and all other hyperparameters between decoders and encoders are equal. To investigate this, we consider two models of comparable size – a small CLM-based decoder and an MLM-based encoder. We randomly initialize small LMs with four decoder or encoder blocks following RoBERTa-base architecture with a language modeling head on top and a newly initialized embedding matrix with the size of RoBERTa-base’s vocabulary. We inherit all other RoBERTa-base hyperparameters and produce a small LM with 68M parameters. We use the RoBERTa-base tokenizer to pre-process the BookCorpus dataset (Zhu et al., 2015) consisting of 74M sentences and use this dataset to pre-train small LMs. We train the small decoder and encoder with CLM and MLM, respectively, with AdamW (Loshchilov and Hutter, 2017) optimizer, a cosine annealing learning rate scheduler, eight gradient accumulation steps, and bfloat16 precision with a batch size of 64 and a learning rate of $2\mathrm{e}{-4}$. We save the model weights immediately after random initialization and then at five equally spaced intervals throughout each epoch (51 checkpoints). The training continues for a fixed number of 10 epochs (roughly 200K steps). After pre-training, we load the weights of each checkpoint, replace the LM head with the SL head for the appropriate task, and fine-tune all parameters for five epochs on a task-specific training set with a batch size of 16. We fine-tune three variants: a pre-trained encoder, a pre-trained decoder with CM during fine-tuning, and a pre-trained decoder without CM. We report averages over five fine-tuning runs, with the last model from each run evaluated on the validation set. More details are provided in Appendix A.2.
@@
The results in Figure 2 reveal that removing the CM on the 68M parameters scale produces no gains. On average, Decoder Unmask performs worse than Decoder Mask. Moreover, encoders struggle to keep up with decoders until around the 20th checkpoint (fourth pre-training epoch), when they start prevailing on all SL tasks except TC. MLM training for over four epochs drastically hurts performance on ACE05. The score RoBERTa-base achieves on ACE05 (cf. Table 2) is close to the best small encoder. The gains from continued MLM training on ACE05 become less important than the overall number of model parameters. This can be explained by the fact that RoBERTa-base and RoBERTa-large were trained for 500k steps (Liu et al., 2019), and RoBERTa-large is drastically better on ACE05.
@@
Figure 2: Validation set micro F1 SL scores of small pre-trained MLM-based encoder and CLM-based decoder models after fine-tuning on SL tasks training data starting from particular LM checkpoint. Decoder Unmask: model pre-trained with CLM and a CM but fine-tuned without the CM. All results are averages over five runs. The shaded area represents the standard deviation.
@@ 6 Conclusion
The landscape of PLMs shifted from encoder to decoder dominance in various NLU tasks. While large decoder models show remarkable performance even without fine-tuning, challenges persist in IE tasks, typically formulated as sequence labeling. In a series of experiments, we showed that LLMs can yield performance competitive with SOTA on sequence labeling IE tasks when the causal mask is removed from specific decoder blocks. Future work should should investigate CM removal for individual layers, analyze task-specific differences in CM removal, and investigate the options for predicting the impact of CM removal without requiring extensive fine-tuning. Comparisons with encoders stress the significance of architecture and scale in SL tasks. A noticeable gap exists between small-scale encoder and decoder models, where decoder models do not benefit from CM removal at a small scale.
7 Limitations
In our experiments, we ensure reliability by averaging performance scores over five runs with different random seeds. Increasing the sample size for averaging would enhance reliability further. However, maintaining fixed learning rates and other hyperparameters across experiments might have led to suboptimal adaptation for sequence labeling tasks. Exploring additional unmasking configurations could provide valuable insights and potential improvements. While numerous open LLMs are available, we focus solely on Llama2 and Mistral, both with 7B parameters. Enhancing prompt templates used for instruction tuning could boost overall instruction tuning performance. Given more computing resources, experimenting with larger models would be feasible. Additionally, our experiments were limited to English-language datasets. Extending the analysis to sequence labeling tasks in other languages and incorporating more diverse datasets could yield further insights.
@@ Appendix A Appendix
A.1 Experimental Setup Details
[…]
For IT experiments, we generate the outputs for evaluation with default generation settings for the Llama2-7B model. To speed up the generation, we decrease the total maximum length of the input instruction prompt combined with newly generated tokens to 1024. The same generation config is used for Mistral-7B.
A.2 Training Small Language Models Details
[…]
A.3 Instruction Tuning Details
@@
Task (dataset) Instruction example for training Instruction example for evaluation NER (CoNLL03) ### Instruction: please extract named entities and their type from the input sentence, all entity types are in options ### Options: person, location, organization, miscellaneous ### Sentence: " What we have to be extremely careful of is how other countries are going to take Germany ’s lead , " Welsh National Farmers ’ Union ( NFU ) chairman John Lloyd Jones said on BBC radio . ### Response: Germany:location;Welsh National Farmers ’ Union:organization;NFU:organization;John Lloyd Jones:person;BBC radio:organization ### Instruction: please extract named entities and their type from the input sentence, all entity types are in options ### Options: person, location, organization, miscellaneous ### Sentence: " What we have to be extremely careful of is how other countries are going to take Germany ’s lead , " Welsh National Farmers ’ Union ( NFU ) chairman John Lloyd Jones said on BBC radio . ### Response: ATE+ATP (Rest14) ### Instruction: please extract aspect terms and their polarity from the input sentence, all polarity types are in options ### Options: positive, negative, neutral, conflict ### Sentence: The lobster sandwich is \$ 24 and although it was good it was not nearly enough to warrant that price . ### Response: lobster sandwich:conflict;price:negative ### Instruction: please extract aspect terms and their polarity from the input sentence, all polarity types are in options ### Options: positive, negative, neutral, conflict ### Sentence: The lobster sandwich is \$ 24 and although it was good it was not nearly enough to warrant that price . ### Response: Trigger Classification (ACE05) ### Instruction: please extract events and their types from the input sentence, all event types are in options ### Options: merge organization, start organization, declare bankruptcy, end organization, grant pardon, extradite, execute, impose fine, conduct trial hearing, issue sentence, file appeal, convict, file lawsuit, release on parole, arrest and send to jail, charge and indict, acquit, participate in protest or demonstration, attack, contact via written or telephone communication, meet, start position, elect, end position, nominate, transfer ownership, transfer money, marry, divorce, be born, die, sustain injury, transport ### Sentence: In his previous letter home , Apache pilot Joe Bruhl did n’t tell his family the full details about his first combat mission . ### Response: tell:contact via written or telephone communication;combat:attack ### Instruction: please extract events and their types from the input sentence, all event types are in options ### Options: merge organization, start organization, declare bankruptcy, end organization, grant pardon, extradite, execute, impose fine, conduct trial hearing, issue sentence, file appeal, convict, file lawsuit, release on parole, arrest and send to jail, charge and indict, acquit, participate in protest or demonstration, attack, contact via written or telephone communication, meet, start position, elect, end position, nominate, transfer ownership, transfer money, marry, divorce, be born, die, sustain injury, transport ### Sentence: In his previous letter home , Apache pilot Joe Bruhl did n’t tell his family the full details about his first combat mission . ### Response: Text Chunking (CoNLL03) ### Instruction: please extract chunks and their type from the input sentence, all chunk types are in options ### Options: noun phrase, verb phrase, prepositional phrase, adverb phrase, subordinated clause, adjective phrase, particles, conjunction phrase, interjection, list marker, unlike coordinated phrase ### Sentence: Rare Hendrix song draft sells for almost \$ 17,000 . ### Response: Rare Hendrix song draft:noun phrase;sells:verb phrase;for:prepositional phrase;almost \$ 17,000:noun phrase ### Instruction: please extract chunks and their type from the input sentence, all chunk types are in options ### Options: noun phrase, verb phrase, prepositional phrase, adverb phrase, subordinated clause, adjective phrase, particles, conjunction phrase, interjection, list marker, unlike coordinated phrase ### Sentence: Rare Hendrix song draft sells for almost \$ 17,000 . ### Response: